Find the interval of convergence for the following power series. 13. (3x – 8) 5n +1 15. (2x – 1) n=1 n=2 ю 14. (2x – 3) (2

Just need 13 14 and 15
Answer #1

§ (32-879 let an= (3x-8) ant1 = (32-83 1+1 n=1 51+1 5nti s(n+1) +1 Snt1 anti 32-8 X Now Im 9200 - lim nt (32-8) 50t6 X snt!